OpenStudy (anonymous):

Sorry I was called out for a moment, abigail2010 or tryeo, if still here, I only have one question, that is why I do not understand using substitution. My problem is, there is 56 houses for sale with two floor plans. I am to use x to represent floor plan one, and y for floor plan two. Then there is 3 times as many homes available with floor plan two, and I am to figure out how many of each floor plans are avaailable using the substitution I am confused and would like someone to explain this to me ste by step if possible cause I want to learn

OpenStudy (anonymous):

56 = x + y right? this is the total number of houses for sale. let x be plan 1 and let y be plan 2. You know that there are 3 times as much homes with 2 than 1, so you know that y = 3x, Substitute 3x into y in the orignal equation, (56 = x +y) and you will get (56 = x+3x) Solve for x and you will get 14.

OpenStudy (anonymous):

when you substitute what exactly are you doing, like why is 56 = x + 3x, what happened to y

OpenStudy (anonymous):

y is equal to 3x.... that was given.. So you can substitute y as 3x.

OpenStudy (anonymous):

It is like when I am counting fruits... In one basket I have 3 apples, and in basket number 2 I have 5 oranges. 3 apples = 3 fruits, and 5 oranges = 5 fruits. So the total number of fruits is 8, or if you wanted to know which types there are, it would be 5 oranges and 3 apples. The variables are interchangeable.

OpenStudy (anonymous):

ok I think I am understanding it a bit more thank you, I am so lost in this math,

OpenStudy (anonymous):

Try it with another problem.. Just understand that If there is an equal sign between two sides, the two sides are equal and therefore you can use them interchangeably.

OpenStudy (anonymous):

This is where I get confused, I do not comprehend writing out the equation in the proper way. Like I can figure it in my head, but it is not the proper way If you understand what I mean. For me, should I first combine the x's and make the equation 56 = 4x

OpenStudy (anonymous):

then divide each side by 4

OpenStudy (anonymous):

Yes that is what you should do... As long as you know the y can be written as 3x.

OpenStudy (anonymous):

You got it.

OpenStudy (anonymous):

Oh yes it just hit me!! Thank you

OpenStudy (anonymous):

Now I am being asked to find the intercepts from part a of this problem which is x + y = 56, then the intercepts from part b, x + 3y =56, then to see where the lines ould intersect if I graph this, the first is to find x right?

OpenStudy (anonymous):

I guess.. I'm not sure I get the question.

OpenStudy (anonymous):

What are the intercepts of this equation, which it is asking the intercepts to graph,